OpenStudy (anonymous):

Find the equation for the slant asymptote for f(x) = (x^2 + 2x)/(7x- 35)

OpenStudy (amistre64):

thats just long division....

OpenStudy (amistre64):

(1/7)x +1 ----------- 7x- 35 | x^2 +2x - x^2 +5x ------------- 7x +0 -7x +35 --------- 35 <- remainder 35 x/7 + 1 + ------- <-- the remainder her gets canceled 7x-35 the oblique asymptote should be: (1/7)x +1

OpenStudy (anonymous):

That makes sense, but I have a question about the long division. Shouldnt the 2x in the first row subtract the 5x underneath it to get -3x instead of add to get 7x?

OpenStudy (amistre64):

No, remember how to do long division with actual numbers... we find the value we are looking for and then "subtract" that value from our problem, like this: 2 <- 2(5) = 10 so we "subtract" that value from "13" ------- 5 |132 -10 ----- 3 same setup, same rules apply.

OpenStudy (amistre64):

since : (1/7)x times (-35) = -5x ; we "subtract" -5x from the answer. +2x -(-5x) = 2x + 5x = 7x